Find the centre and radius of the sphere passing through (1,0,0),(0,1,0),(0,0,1) and (1/√3,1/√3,1/√3)?
1
Expert's answer
2020-06-01T18:10:27-0400
"(x-x_c)^2+(y-y_c)^2+(z-z_c)^2=R^2"

"(1,0,0): (1-x_c)^2+(0-y_c)^2+(0-z_c)^2=R^2"

"(0,1,0): (0-x_c)^2+(1-y_c)^2+(0-z_c)^2=R^2"

"(0,0,1): (0-x_c)^2+(0-y_c)^2+(1-z_c)^2=R^2"



"x_c^2-2x_c+1+y_c^2+z_c^2=R^2""x_c^2+y_c^2-2y_c+1+z_c^2=R^2""x_c^2+y_c^2+z_c^2-2z_c+1=R^2"


"x_c=y_c=z_c""3x_c^2-2x_c+1=R^2"

"(1/\sqrt{3}, 1/\sqrt{3}, 1/\sqrt{3}):"


"({1\over \sqrt{3}}-x_c)^2+({1\over \sqrt{3}}-y_c)^2+({1\over \sqrt{3}}-z_c)^2=R^2"

Then


"3({1\over \sqrt{3}}-x_c)^2=3x_c^2-2x_c+1"

"1-2\sqrt{3}x_c+3x_c^2=3x_c^2-2x_c+1"

"x_c={1\over \sqrt{3}}={\sqrt{3}\over 3}=y_c=z_c"

"R^2=3({1\over \sqrt{3}})^2-2({1\over \sqrt{3}})+1=2-{2\over \sqrt{3}}"

"C({\sqrt{3}\over 3}, {\sqrt{3}\over 3}, {\sqrt{3}\over 3}), R=\sqrt{2-{2\over \sqrt{3}}}"